# Keeping in Place After the Storm—Emergency Assistance and Evictions

### v1

@@ Abstract
We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided. We find a 16% increase in overall evictions after hurricane natural disaster events that can be decomposed to a 36% increase for those that do not receive FEMA rental assitance and only a 11% for hurricane events that do receive FEMA aid. Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes remain stable and result in a decrease in defaults by 19% in payday loans during hurricane events in locations that do receive FEMA aid. This effect largely reverses in areas that do not receive FEMA aid, where transaction volumes drop by 12% and default rates remain similar relative to the baseline. Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
@@ 1 Introduction
The loss of structures and in particularly housing in the wake of natural disasters is often among the notable and lasting images of large natural disasters. Recent catastrophes such as Hurricane Katrina in 2005 and Hurricane Harvey in 2017, continue to attract attention because of the their large-scale visible impacts, notably the destruction of buildings and neighborhoods, the resultant displacement and the post-mortem evaluation of government responses. While the response to such events may yield silver linings in the ensuing learned adaptation, at the micro level, the loss born by individuals can be large and has been documented in various forms in the literature.
Historical instances in which neighborhoods that have been rebuilt do not necessarily resemble their former composition as a result of the immediate loss of housing in such events. The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States. To contrast, in our work, we study instances of assistance in preserving housing access. That is, where federal assistance exists to enable households to maintain their ability to make housing payments when potential disruptions directly related to the storm event affect the ability to make such payments.
@@ 1.1 Literature Review
Furthermore, the consequences of hurricane events specifically has more generally received attention in the literature, (Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances. (Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions. (Del Valle et al., 2024) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster. They identify shifts in financial behavior, including increased reliance on savings and external credit. Similarly, (Collier et al., 2024) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions. Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
@@ 2.1.4 FEMA Data
Disaster Name Registered for FEMA Aid Approved for FEMA Aid Total IHP Amount Disbursed Total Rental Assistance Disbursed Total IHP Amount per Approved Applicant
Tropical Storm Bonnie and Hurricane Charley 50,136 28,945 70,702,271 26,273,487 2,443
Hurricane Frances 114,540 66,784 156,876,161 46,208,593 2,349
Hurricane Ivan 35,961 21,203 52,330,846 11,684,323 2,468
Hurricane Jeanne 89,278 57075 153,826,507 39,083,876 2,695
Hurricane Dennis 9,131 4238 4,018,903 1,006,428 948
Hurricane Wilma 150,501 57,542 92,416,813 23,939,083 1,606
Tropical Storm Fay 5,463 1622 3,395,589 2,098,470 2,093
Tropical Storm Debby 4,455 1,752 4,555,721 2,771,489 2,600
Hurricane Hermine 1,643 585 1,772,365 797,293 3,030
Hurricane Irma 1,420,062 444,735 495,234,951 304,094,846 1,114
@@ 2.2.2 Event Study Framework
As in (Doruk Cengiz and Zipperer, 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously. This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
@@ 2.2.3 Difference-in-Differences Framework
Finally, to shed light on the role of FEMA aid, we conduct an interacted difference-in-differences comparison for hurricane events that are declared FEMA and resulted in the availability of rental assistance (or were declared FEMA but did not result in any rental assistance). For this approach, the control group is the same as the event based matched control group in the earlier difference-in-difference; in contrast, we divide the treated group into two groups. The non-FEMA treated group contains ZIP codes that were impacted by hurricanes or tropical storms but didn’t receive FEMA aid, and the treated FEMA group is ZIP codes impacted by hurricanes and tropical storms that received FEMA aid.
@@ 3.1.1 Hurricanes
Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. Coefficients for Treated and Post are not reported, as they are absorbed by fixed effects. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. The coefficient of 0.525 for evictions represents an approximate 16.2% increase relative to the treated group’s pre-event average. Similarly, the 0.928 increase in filings corresponds to about a 17.9% rise from the treated group’s pre-event average. Standard errors clustered at the ZIP code level are reported in square brackets. ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
@@
Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
[0.172] [0.235]
@@
Eviction Rate 3.665 3.238
Filing Rate 5.592 5.182
@@
We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes. Our results show a much larger bifurcation in the resultant filings and evictions in the period following the hurricane event. Both the event-study and difference-in-differences regression results illuminate the increased rate of filings and evictions for hurricane events where FEMA assistance is not made available. In particular, we find an increased monthly average of 1.115 and 1.543, for evictions and filings per month, respectively Table 6. These represent a 35.7% and 29.8% increase, respectively, for hurricanes that did not receive FEMA aid. Comparatively, the availability of FEMA rental assistance resulted in a 11.2% increase over the baseline in evictions. The absence of FEMA assistance resulted in three times the increase in the rate of evictions and two times the increase in the rate of filings. This also highlights that there was not only a greater increase in the number of evictions but a comparatively greater increase in the rate of successful evictions relative to filings, given the greater increase of evictions to filings.
@@
Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction rates at the ZIP code level. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the monthly eviction rate. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level. The estimates suggest a marked post-event increase in eviction rates in non-recipient ZIP codes, whereas areas receiving aid experienced relatively muted changes, indicating a potential mitigating role of disaster assistance.
@@
Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly filing rates at the ZIP code level. The event window covers five months prior to and six months following the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays results for ZIP codes that received FEMA aid, while Panel (b) shows results for ZIP codes that did not receive FEMA aid. The dependent variable is the monthly filing rate. Vertical lines indicate 95% confidence intervals, calculated using standard errors clustered at the ZIP code level. The findings reveal a substantial post-event increase in filing rates in non-recipient areas, whereas the increase is more muted in aid-recipient ZIP codes—suggesting that FEMA assistance may play a mitigating role in the aftermath of natural disasters.
Table 6: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, estimated separately for FEMA and non-FEMA designated areas. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For FEMA areas, the coefficient of 0.363 for evictions represents an approximate 10.6% increase relative to the treated group’s pre-event average (3.427), while the 0.769 increase in filings corresponds to about a 14.3% rise from the treated group’s baseline (5.392). In non-FEMA areas, the 1.155 increase in evictions indicates a roughly 50.1% rise from the treated pre-event average (2.307), and the 1.543 increase in filings represents a 37.2% increase from the baseline (4.143).
Dependent Variable Eviction (FEMA) Filing (FEMA) Eviction (No-FEMA) Filing (No-FEMA)
Period Monthly Monthly Monthly Monthly
Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.363 0.769∗∗∗ 1.155∗∗∗ 1.543∗∗∗
[0.188] [0.251] [0.353] [0.510]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 11,097 11,097 4,544 4,544
Variable Control Group Treated Group
FEMA No FEMA FEMA No FEMA
Eviction Rate 3.578 3.869 3.427 2.307
Filing Rate 5.298 6.273 5.392 4.143
Table 8: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. The specification includes an interaction between Treated, Post, and a FEMA designation indicator, allowing the treatment effect to vary by FEMA status. The estimation window spans 5 months before to 6 months after the hurricane event. The coefficient on Treated×$\times$×Post indicates the average treatment effect in non-FEMA areas: a 1.155 increase in monthly evictions (approximately 50.1% relative to the treated pre-event mean of 2.307) and a 1.543 increase in filings (a 37.2% increase from the baseline of 4.143). The negative Treated×$\times$×Post×$\times$×FEMA interaction suggests that the treatment effect was significantly smaller in FEMA-designated areas: 0.792 fewer evictions and 0.775 fewer filings, relative to non-FEMA treated ZIP codes.
@@
Treated×$\times$×Post 1.155∗∗∗ 1.543∗∗∗
[0.353] [0.510]
Treated×$\times$×Post×$\times$×FEMA -0.792∗∗ -0.775
[0.388] [0.548]
@@
Observations 15,641 15,641
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
@@ 3.1.2 Hurricanes and Tropical Storms
Eviction Rate 3.937 3.113
Filing Rate 5.808 5.033
We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included. Although, the comparative magnitudes decreases and and the difference due to availability of FEMA assistance also is not statistically different than zero. We show that this is largely due to tropical storms seeing far lower post-storm eviction rates. Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events. Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the absence of FEMA assistance.
@@
Figure 8: The figure provides the estimated coefficients from an event study examining the effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level. The results indicate a notable post-event increase in both eviction and filing rates, highlighting the disruptive impact of storm exposure on evictions.
@@
Table 10: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storms and another excluding hurricane events. Coefficients for Treated and Post are not reported, as they are absorbed by event-specific ZIP code and time fixed effects. The estimation window spans from 5 months before to 6 months after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.371 evictions and 0.633 filings in treated ZIP codes. These correspond to approximately 11.9% and 12.6% increases relative to the treated group’s pre-event averages of 3.113 and 5.033, respectively. When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
@@
Treated×$\times$×Post 0.371∗∗∗ 0.633∗∗∗ 0.128 0.171
[0.138] [0.186] [0.218] [0.283]
@@
Observations 48,076 48,076 32,435 32,435
Table 11: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event. The coefficient of 0.525 for evictions and 0.928 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 16.9% and 18.4% increases relative to the pre-event averages, respectively. The interaction term Treated×$\times$×Post×$\times$×Storm represents the differential impact of storms. For evictions, the coefficient is -0.398, not statistically significant, while for filings, the coefficient is -0.757∗∗, indicating a statistically significant reduction in filings in treated ZIP codes exposed to tropical storms.
@@
Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
[0.172] [0.235]
Treated×$\times$×Post×$\times$×Storm -0.398 -0.757∗∗
[0.271] [0.360]

### v2

@@ Abstract
We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided. We find an approximate 16.5% increase in overall evictions after hurricane natural disaster events that increases to 19.7% when excluding ZIP codes that receive FEMA rental assistance. Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes and defaults decrease during hurricane events in locations that do receive FEMA aid. This effect largely reverses in areas that do not receive FEMA aid, where the magnitude of transaction volumes drop by less and default rates remain similar relative to the baseline. Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
@@ 1 Introduction
The loss of structures and in particularly housing is often among the notable and lasting images of in the wake of natural disasters. The resulting financial loss and displacement highlight critical questions over both the individual and government role in pre-disaster mitigation and post-disaster response, in particular through their shortcomings. At the micro level, the loss born by households can be large and has been documented in various forms in the literature.
Historically, the destruction of housing in such events has often led to rebuilt neighborhoods that differ significantly from their original form and community composition. The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States. To contrast, in our work, we study instances of assistance in preserving housing access. That is, we examine the existence of federal assistance households to maintain their ability to make housing payments due to potential disruptions and damages directly related to the natural disaster.
@@ 1.1 Literature Review
Furthermore, the consequences of hurricane events has notably received attention in the literature. (Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances. (Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions. (Del Valle et al., 2024) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster. They identify shifts in financial behavior, including increased reliance on savings and external credit. Similarly, (Collier et al., 2024) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions. Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
@@ 2.1.4 FEMA Data
Disaster Name
Registered for
FEMA Aid
Approved for
FEMA Aid
Total Rental
Assistance Disbursed (\$)
Average Rental Assistance
per Recipient (\$)
Tropical Storm Bonnie and
Hurricane Charley
50,136 28,945 26,273,487 912
Hurricane Frances 114,540 66,784 46,208,593 695
Hurricane Ivan 35,961 21,203 11,684,323 554
Hurricane Jeanne 89,278 57075 39,083,876 687
Hurricane Dennis 9,131 4238 1,006,428 242
Hurricane Wilma 150,501 57,542 23,939,083 418
Tropical Storm Fay 5,463 1,622 2,098,470 1,342
Tropical Storm Debby 4,455 1,752 2,771,489 1,638
Hurricane Hermine 1,643 585 797,293 1,442
Hurricane Irma 1,420,062 444,735 304,094,846 685
@@ 2.2.2 Event Study Framework
As in (Cengiz et al., 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously. This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
@@ 2.2.3 Difference-in-Differences Framework
Finally, to shed light on the role of FEMA aid, we conduct an interacted event study comparison on the treated ZIP codes for hurricane events and resulted in ZIP codes having the availability of rental assistance. In addition we also conducted a difference-in-differences analysis while excluding treated units that received any FEMA assistance, therefore only on the non-FEMA receiving ZIP codes. For the difference-in-difference approach, the control group is the same as the event based matched control group in the earlier difference-in-difference; in contrast, for the FEMA event-study on the treated, we separated ZIP codes into those that received FEMA assistance and those do not within the same hurricane event.
@@ 3.1.1 Hurricanes
Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. the coefficient of 0.331 for evictions represents an approximate 16.5% increase relative to the treated group’s pre-event average. Similarly, the 0.428 increase in filings corresponds to about a 13.3% rise from the treated group’s pre-event average. Standard errors clustered at the ZIP code level are reported in square brackets. ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
@@
Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
[0.089] [0.110]
@@
Eviction Rate 2.403 2.002
Filing Rate 3.601 3.203
@@
We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes. Our results indicate a possible bifurcation in the resultant filings and evictions in the period following the hurricane event. Both the event-study and difference-in-differences regression results illuminate the decrease rate of filings and evictions for hurricane events where FEMA assistance is made available. In particular, we firstly find a decrease in the monthly average of 0.418 and 0.445, for evictions and filings per month, respectively Table 6 when restricting to the event study just within treated units. These represent a 20.9% and 13.9% decrease, respectively, for hurricanes that did not receive FEMA aid. Although both results are not statistically significant, we also show a larger estimate for evictions and filings when excluding ZIP codes that included FEMA assistance in the difference-in-differences results relative to the baseline. Comparatively, evictions increased by a statistically significant 0.392 when excluding FEMA receiving ZIP codes relative to the control group, a larger coefficient than 0.331 result when including all ZIP codes. Similarly, filings increased by a statistically significant 0.502 when excluding FEMA assistance, larger than the 0.428 coefficient when including all ZIP codes.
@@
Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level on the treated. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents the interacted monthly eviction estimates for ZIP codes that received FEMA aid versus those that did not for the treated group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that received FEMA aid. Standard errors are clustered at the ZIP code level.
@@
Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level with the entire set of control ZIP codes but excluding the treated ZIP codes that do not receive FEMA assistance. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents the interacted monthly eviction estimates for ZIP codes that do not receive FEMA assistance versus the control group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that do not receive FEMA assistance versus the control group. Standard errors are clustered at the ZIP code level.
@@
FEMA×$\times$×Post -0.418 -0.445
[0.291] [0.303]
@@
Observations 2729 2729
Table 6: Results for monthly eviction and filing outcomes at the ZIP code level on the treated group only, separated into ZIP codes that received FEMA assistance and those did not. this provides within storm event estimate comparisons for all ZIP codes that were within the storm path. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For FEMA receiving ZIP codes, the coefficient of -0.418, for evictions, while not statistically significant, represents an approximate 20.9% decrease relative to the treated group’s pre-event average (2.002), while the -0.445 decrease in filings corresponds to about a 13.9% decrease from the treated group’s baseline (3.203).
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.392∗∗∗ 0.502∗∗∗
[0.092] [0.116]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 16647 16647
Table 7: Difference-in-difference results for monthly eviction and filing outcomes at the ZIP code level while excluding treated ZIP codes that received FEMA assistance. This provides a comparison with control group ZIP codes that for the same time period were not in the storm path. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For non-FEMA receiving ZIP codes, the statistically significant coefficient of 0.392, for evictions, is larger than the 0.331 coefficient for all treated ZIP codes, represents an approximate 19.6% increase relative to the treated group’s pre-event average (2.002), while the 0.502 increase in filings, larger than the 0.428 coefficient for all treated ZIP codes, corresponds to about a 15.6% increase from the treated group’s baseline (3.203).
@@ 3.1.2 Hurricanes and Tropical Storms
Eviction Rate 2.664 2.423
Filing Rate 3.950 3.829
We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included. Although, the comparative magnitudes decreases, we show that this is largely due to tropical storms seeing far lower post-storm eviction rates. Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events. Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance.
@@
Figure 8: The figure provides the estimated coefficients from an event study examining the effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
@@
Table 9: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storms and another excluding hurricane events. The estimation window spans from 5 months before to 6 months after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.209 evictions and 0.336 filings in treated ZIP codes. These correspond to approximately 8.6% and 8.7% increases relative to the treated group’s pre-event averages of 2.423 and 3.829, respectively. When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
@@
Treated×$\times$×Post 0.209∗∗∗ 0.336∗∗∗ 0.103 0.255
[0.070] [0.125] [0.105] [0.200]
@@
Observations 62209 62209 43708 43708
Table 10: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event. The coefficient of 0.331 for evictions and 0.428 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 13.6% and 11.1% increases relative to the pre-event averages of 2.423 and 3.829, respectively. The triple interaction Treated×$\times$×Post×$\times$×Storm is negative and marginally significant (-0.228, significant at the 10%) for Eviction, suggesting that the treatment effect after tropical storms is somewhat lower than after hurricanes. In other words, tropical storms appear to have a weaker impact on evictions compared to hurricanes. However, for eviction filings, the triple interaction (-0.173) is negative but not statistically significant.
@@
Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
[0.089] [0.110]
Treated×$\times$×Post×$\times$×Storm -0.228 -0.173
[0.137] [0.212]